# Dynamical discontinuities in repeated weak measurements revealed by complex weak values

### v1

@@ Abstract
Critical phenomena reveal universal behavior in complex systems, and uncovering analogous effects in quantum weak measurement protocols with post-selection provides new insight into how measurement backaction can shape quantum dynamics. This work investigates dynamical discontinuities that arise when the post-selected polar angle is used as a control parameter. The system evolves under repeated applications of a weak measurement protocol with post-selection, in which the meter state is retained after each iteration while the system is renewed. The emergence of these discontinuities is shown to be determined by the structure of the weak value: when the weak value has a nonzero imaginary component, a discontinuity appears in the expectation value of meter observables precisely at the point where the imaginary part of the weak value vanishes as a function of the post-selection polar angle. In contrast, no discontinuities occur when the weak value remains real for all post-selection angles. The phenomenon originates from the eigenstructure of the protocol’s Kraus operator, with the stability of fixed points changing at the critical point where the discontinuity arises. Remarkably, the associated critical exponent is $1$, independent of system parameters. These results open new perspectives for engineering non-analytic behavior in measurement-based quantum control and for probing criticality in post-selected quantum dynamics using weak measurements with weak values.
@@ I Introduction
Weak measurements with post-selection, first introduced by Aharonov and colleagues in [aharonov1988result], reveal a shift in the meter proportional to the so-called weak value, which can lie outside the eigenvalue spectrum of the measured observable. These anomalous weak values have proven useful both in foundational studies of quantum theory [aharonov2002revisiting, matzkin2012observing, dressel2015weak, ferraz2022geometrical, ferraz2023revisiting, ferraz2025exploring] and in enhancing sensitivity to weak signals in precision measurement and sensing applications [jordan2014technical, dixon2009ultrasensitive, luo2017precision, ferraz2024relevance]. Furthermore, weak values are closely connected to the Kirkwood–Dirac quasiprobability distribution, which exhibits nonclassical features whenever weak values are anomalous. In addition, anomalous weak values serve as evidence of contextuality in quantum mechanics [arvidsson2024properties, pusey2014anomalous, kunjwal2019anomalous].
The study of non-equilibrium quantum dynamics has revealed a rich landscape of critical phenomena that extend the concept of phase transitions beyond equilibrium physics. In addition to traditional equilibrium transitions driven by temperature or external fields [sachdev1999quantum, Hertz1976, millis1993effect, vojta2003quantum], non-equilibrium settings exhibit sharp, non-analytic behavior in time-evolving observables [heyl2013dynamical, vzunkovivc2016dynamical] and entanglement measures [de2021entanglement, kawabata2023entanglement]. Dynamical phase transitions, characterized by non-analyticities in quantities such as the Loschmidt echo following a quantum quench, have been widely explored in spin chains, cold-atom systems, and trapped-ion simulators, uncovering deep connections between dynamical criticality, equilibrium transitions, and excited-state quantum phase transitions [heyl2013dynamical, zvyagin2016dynamical, jurcevic2017direct]. In parallel, measurement-induced phase transitions have highlighted the fundamental competition between unitary entanglement generation and measurement-induced decoherence, leading to a transition between volume-law and area-law entangled phases in many-body systems [skinner2019measurement, block2022measurement, li2023cross]. These measurement-driven transitions exhibit universal scaling behavior reminiscent of classical percolation and have been observed in both numerical simulations and quantum experiments [noel2022measurement, koh2023measurement]. Remarkably, signatures of criticality and non-analytic behavior are not restricted to the thermodynamic limit: even minimal systems, such as two-level and Rabi-type models, can display dynamical discontinuities when system parameters are tuned across critical regimes, suggesting that essential features of quantum criticality persist in few-body settings [puebla2020finite, hamazaki2021exceptional, di2024environment].
Although weak measurements with post-selection have been associated with mathematical features related to quantum criticality, such as non-normality [ferraz2023revisiting], they have not yet been connected to critical phenomena, nor has the weak value been studied as a potential indicator of such behavior. This work investigates these aspects within a repeated weak measurement protocol.
This work investigates discontinuities, analogous to phase transitions, that emerge when the post-selection polar angle serves as the control parameter. Such discontinuities manifest in the expectation value of a meter observable after repeatedly applying the weak measurement protocol, as defined in [aharonov1988result], to the meter state. Analysis of the protocol’s structure reveals the conditions under which discontinuities emerge. Importantly, my focus is not on the thermodynamic limit of many-body systems but rather on discontinuities within small systems. As a possible extension, one could explore whether enlarging the system Hilbert space enables genuine phase transitions in the thermodynamic sense.
@@
The study examines how discontinuities emerge in the expectation value of meter observables during repeated weak measurement interactions. In each iteration, the system and meter interact via a weak unitary coupling, characterized by small interaction strength and short duration, after which the system is post-selected. The system is then discarded, while the meter is retained to interact with a new system initialized in the same initial state, following the same procedure. After $N$ iterations, the expectation value of an observable (e.g., a Pauli operator $\hat{\sigma}$ for a two-level system) is evaluated as a function of the post-selection polar angle. Discontinuities arise precisely when the imaginary part of the weak value vanishes, provided that (i) the weak value is not purely real for all post-selection angles and (ii) the Kraus operator governing the meter’s evolution has nondegenerate eigenvalues.
The study of quantum dynamics under repeated measurements has a long history, from the rigorous analyses of quantum nondemolition protocols [Bauer2011, Bauer2012] to the exploration of measurement-induced phase transitions in many-body systems [Szyniszewski2019, Dubey2021]. In quantum nondemolition protocols and weak-measurement settings, repeated probe interactions lead to convergence toward stable fixed points or steady states, with convergence rates governed by the eigenstructure of the corresponding Kraus maps. However, these processes are typically continuous in their control parameters and do not exhibit sharp dynamical discontinuities. In contrast, many-body measurement-induced transitions [Szyniszewski2019] show non-analytic behavior in entanglement measures as a function of measurement strength, although the mechanism arises from competition between unitary and non-unitary dynamics. More recently, repeated weak measurements of non-conserved observables have been shown to yield partial collapse and emergent steady-state structures [Roux2024], but again without the appearance of discontinuous expectation values. This work differs by demonstrating that discontinuities can emerge even in a minimal two-particle setting, when the imaginary component of the weak value crosses zero, producing an exchange of stability bifurcation in the dynamics of the associated Kraus map.
Extending the phase-transition analogy, the relaxation dynamics reveal a critical exponent of 1, independent of system parameters, highlighting a form of universality. Nevertheless, it should be emphasized that these discontinuities do not constitute true phase transitions, as no thermodynamic limit is involved. However, by extending the meter dimension, one could potentially approach such a limit.
@@ II Protocol presentation
This section introduces the protocol that gives rise to discontinuities and stability bifurcations, and examines the associated Kraus operator. Several qubits serve as the “systems” in a standard weak measurement protocol with post-selection, while an $N$-level system acts as the meter, the system in which discontinuities are induced.
@@
The system and meter interact weakly through a unitary operator $\hat{U}$ which, assuming the interaction strength $g$ and the time $t$ are sufficiently small, can be expanded to first order in $g$ as
@@
is the weak value of the observable $\hat{O}_{S}$.
@@
The weak measurement protocol to reach the discontinuity (Fig. 1), consisting of pre-selection, a weak unitary interaction, and post-selection, is applied to the meter, with the entire process represented by the Kraus operator $\hat{K}$. After each full application of the weak measurement protocol, the system is discarded and replaced with a new one for the next iteration, while the meter is preserved throughout all $n$ applications.
@@ III Fixed points and their stability
$\hat{O}_{A}|a_{j}\rangle=\lambda_{j}|a_{j}\rangle,\qquad\lambda_{j}\in\mathbb{R},$ (9)
where $j$ runs from 1 to $N$, the dimension of the Hilbert space of the operator $\hat{O}_{A}$, and $\lambda_{j}$ are its eigenvalues.
@@
$|c_{j}^{(n)}|^{2}=\frac{|c_{j}^{(0)}|^{2}\,\left|1+2gt\,\text{Im}(O_{S,w})\,\lambda_{j}\right|^{2n}}{\sum_{k}|c_{k}^{(0)}|^{2}\left|1+2gt\,\text{Im}(O_{S,w})\,\lambda_{k}\right|^{2n}},$ (11)
@@
As $n\to\infty$, the system approaches a long-time limit characterized by a stable fixed point, which is the state toward which all other states converge, except for those corresponding to unstable fixed points. This stable state is the eigenvector of the Kraus operator associated with the largest value of $|1+\alpha\lambda_{j}|$, assuming that the Kraus operator spectrum is non-degenerate. The remaining eigenvectors correspond to unstable fixed points.
@@
  • If $\text{Im}(O_{S,w})>0$, the stable state is the eigenvector with the largest eigenvalue $\lambda_{+}$.
  • If $\text{Im}(O_{S,w})<0$, the stable state is the eigenvector with the smallest eigenvalue $\lambda_{-}$.
@@ IV Qubit-qubit example
Figure 2: Three distinct discontinuities characterize the dependence of the expectation value of $\hat{\sigma}_{x}$ on the post-selected angle $\phi$, emphasizing the nontrivial structure that arises for different repetition numbers $n$. In this plot $\alpha=\frac{\pi}{7}$, $\theta=\frac{\pi}{4}$ and $\gamma=0.001$, while the initial meter state is along the $z$ axis.
@@
Figure 3: The imaginary part of the weak value vanishes when the absolute value of the two eigenvalues of the Kraus operator coincide. The plot shows the eigenvalues of the Kraus operator $\hat{K}$, $|\lambda_{+}|$ and $|\lambda_{-}|$, as functions of the post-selected angle $\phi$, together with their difference $|\lambda_{+}|-|\lambda_{-}|$ and the imaginary part of the weak value $\mathrm{Im}(O_{s,w})$. In this plot $\alpha=\frac{\pi}{7}$, $\theta=\frac{\pi}{4}$ and $\gamma=0.1$, while the initial meter state is along the $z$ axis.
As predicted analytically, the eigenvalues have equal absolute values whenever the imaginary part of the weak value vanishes (Fig. 3).
@@
$r_{x,n}=\text{Tr}\left[\hat{\sigma}_{x}\hat{\rho}_{A}^{n}\right]=\frac{ar_{x,n-1}+b+c}{a+\left(b+c\right)r_{x,n-1}}$ (19)
$r_{y,n}=\text{Tr}\left[\hat{\sigma}_{y}\hat{\rho}_{A}^{n}\right]=\frac{ar_{y,n-1}+ir_{z,n-1}\left(b-c\right)}{a+\left(b+c\right)r_{x,n-1}}$
$r_{z,n}=\text{Tr}\left[\hat{\sigma}_{z}\hat{\rho}_{A}^{n}\right]=\frac{ar_{z,n-1}-ir_{y,n-1}\left(b-c\right)}{a+\left(b+c\right)r_{x,n-1}},$
where $a=|c\left(\gamma\right)|^{2}+|d\left(\gamma\right)|^{2}$, $b=c\left(\gamma\right)d^{*}$, $c=c^{*}\left(\gamma\right)d\left(\gamma\right)$.
@@
For values of $\phi$ such that $\phi\ll\frac{\pi}{2}$, two fixed points emerge, corresponding to the eigenvectors of $\hat{\sigma}_{x}$, the operator acting on the meter in Eq. 15. The fixed point at x=1x=-1 is stable, while the one at $x=1$ is unstable. Even if the state begins close to $x=1$, the dynamics will eventually drive it toward the stable fixed point at x=1x=-1 (Fig. 5). The dynamics will remain at $x=1$ only when the system is initially exactly at that point. Conversely, for $\frac{\pi}{2}\ll\phi\ll\pi$, the stability of the fixed points is reversed: $x=1$ becomes stable and x=1x=-1 unstable. After passing $\phi=\frac{\pi}{2}$, the stable and unstable fixed points effectively swap roles. Finally, in the near-unitary regime, when the dynamics are close to but not exactly unitary, the trajectories cycle on the Bloch sphere before eventually collapsing onto the stable fixed point (Fig. 6).

### v2

@@ Abstract
This work demonstrates that repeated weak measurements together with post-selection can produce sharp dynamical discontinuities in meter observables, even in minimal quantum systems. The discontinuous behavior is governed by the polar angle of the post selected state, which serves as a continuous control parameter. As this angle is varied, the expectation value of the meter observables changes abruptly at the point where the imaginary part of the associated weak value, a complex quantity that arises in weak measurements with post-selection, becomes zero. Such non-analytic behavior emerges only when the weak value is genuinely complex for a range of post-selection angles. If the weak value remains purely real for all angles, the dynamics remain smooth. The discontinuity originates from an exchange of stability between fixed points of the non-unitary Kraus operator governing the meter’s evolution. Remarkably, despite the absence of a thermodynamic limit, the relaxation time in the vicinity of the discontinuity exhibit universal critical behavior characterized by a critical exponent equal to $1$, independent of system parameters. These results establish the weak value as a tunable control parameter capable of inducing non-analytic dynamical responses and reshaping the stability structure of measurement-induced quantum dynamics.
@@ I Introduction
Weak measurements with post-selection, introduced by Aharonov and colleagues [1], consist of three steps. First, the system is prepared in a known initial state, a procedure referred to as pre-selection. Second, it undergoes a weak and short interaction with a measurement device, described by a unitary operator with a small coupling strength. Finally, a specific final state of the system is post-selected through a projective measurement followed by filtering. In this setting, the meter shift is proportional to the weak value of the observable, a complex quantity that can be anomalous, meaning it may lie outside the range of the observable’s eigenvalues. For example, it can take values larger or smaller than the eigenvalue spectrum or have a nonzero imaginary part.
The physical interpretation of complex weak values has been explored in the literature. Jozsa analyzed how both the real and imaginary parts of the weak value manifest in measurable shifts of the pointer’s position and momentum in a weak measurement context [25]. Hofmann showed that the complex phases that arise in weak conditional probabilities describe the dynamical response of a system to small unitary transformations generated by the measured observable [22, 23]. These studies connect the complex nature of weak values with concrete measurement outcomes and underlying dynamical properties of quantum systems.
@@
Anomalous weak values, whose values are outside the observable spectrum, have proven useful both in foundational studies of quantum theory [2, 35, 10, 14, 16, 13] and in enhancing sensitivity to weak signals in precision measurement and sensing applications [24, 9, 34, 15]. Furthermore, weak values are closely connected to the Kirkwood–Dirac quasiprobability distribution, which exhibits nonclassical features whenever weak values are anomalous. Anomalous weak values have been associated with quantum contextuality, as well as with counterfactual and coherence effects in pre- and post-selected scenarios [3, 42, 31, 19, 18].
The study of non-equilibrium quantum dynamics has uncovered a rich variety of critical phenomena that extend the concept of phase transitions outside the realm of equilibrium physics. While traditional equilibrium transitions are driven by temperature or external fields, such as magnetic fields [44, 20, 36, 47], non-equilibrium quantum systems can exhibit purely quantum phase transitions with no classical analogue. These transitions may manifest in the dynamics of time-evolving observables [21, 48] or in measures of quantum entanglement [7, 27].
Dynamical phase transitions arise in quantum systems driven out of equilibrium, typically after a sudden change in a system parameter, called a quantum quench. Unlike conventional phase transitions, which involve abrupt changes in thermodynamic quantities as external control parameters are varied, dynamical phase transitions are marked by non-analytic behavior in time-dependent observables. A common example is the Loschmidt echo, which quantifies the overlap between the initial state and its time-evolved state; sharp kinks or sudden changes in this quantity indicate a qualitative change in the system’s dynamics. Such transitions have been observed in spin chains, cold-atom systems, and trapped-ion simulators [21, 49, 26].
Measurement-induced phase transitions emerge from the competition between unitary entanglement generation and measurement-induced decoherence. While unitary dynamics tend to spread entanglement, repeated measurements suppress it, leading to a phase transition between highly entangled and weakly entangled states with universal scaling behavior [45, 6, 33]. These transitions have been observed in both numerical simulations and quantum experiments [38, 30], and in many-body systems they exhibit genuine non-analytic behavior, most prominently in entanglement measures, as the measurement rate or strength is varied [46].
Remarkably, signatures of criticality and non-analytic behavior are not restricted to the thermodynamic limit. Even small systems, such as two-level or Rabi-type models, can display sudden dynamical changes when system parameters are tuned across critical regimes. This indicates that essential features of quantum criticality, including abrupt transitions in observables and entanglement, persist in few-body settings [41, 17, 8].
The study of quantum dynamics under repeated measurements has a long history, ranging from rigorous analysis of quantum non-demolition protocols [4, 5] to the investigation of measurement-induced phase transitions in many-body systems [46, 11]. In quantum non-demolition and weak-measurement settings, a system repeatedly interacts with probe degrees of freedom, and each measurement updates the state according to a completely positive map (a Kraus map) describing the measurement back-action. Iterating this map typically drives the system toward a stable fixed point or steady state, with the rate of convergence determined by the eigenvalues of the map. Importantly, in such scenarios the dependence on control parameters, such as measurement strength, is generally smooth, leading to gradual relaxation rather than abrupt changes in observable expectation values. More recently, repeated weak measurements of non-conserved observables have been shown to induce partial state collapse and emergent steady-state structures [43], although the evolution of observables remains continuous as system parameters are tuned.
This work investigates the dynamics of quantum systems under repeated weak measurements with post-selection, where sharp discontinuities can appear in the expectation values of meter observables. The protocol shares certain features with measurement-induced and dynamical phase transitions, yet also differs in important ways. Unlike these phase-transition protocols, the focus here is on small systems. In contrast to measurement-induced transitions, which are usually associated with entanglement dynamics, this study focuses on expectation values, despite the presence of measurements. There is also an analogy with dynamical phase transitions, where post-selection plays a role similar to a quench, although the control parameter is not time. While weak measurements with post-selection have been linked to mathematical features reminiscent of quantum criticality, such as non-normality [16], they have not previously been connected to critical phenomena, nor has the weak value been systematically studied as an indicator of such behavior.
@@
The protocol proceeds as follows. In each iteration, the system is initialized in a chosen state (pre-selection) and interacts weakly with a meter via a short, low-strength unitary coupling. Afterward, the system is post-selected in a desired state and discarded, while the meter is retained to interact with a freshly initialized system. Repeating this procedure $N$ times allows evaluation of the expectation value of a meter observable, for example, a Pauli operator $\hat{\sigma}$ in a two-level system, as a function of the post-selection polar angle. Discontinuities arise precisely when the imaginary part of the weak value vanishes, provided that (i) the weak value is not purely real for all post-selection angles, and (ii) the Kraus operator governing the meter evolution has non-degenerate eigenvalues.
The discontinuity is triggered when the imaginary component of the weak value crosses zero, inducing an exchange-of-stability bifurcation in the Kraus map: two dynamical fixed points interchange their stability, producing an abrupt change in the long-time behavior of the meter observable.
Extending the analogy with phase transitions, the relaxation dynamics near the discontinuity exhibit power-law scaling with a critical exponent of 1. Specifically, as the control parameter, the post-selection polar angle, approaches its critical value where the imaginary part of the weak value vanishes, the characteristic relaxation time, defined as the number of measurement iterations needed for the meter observable to reach its steady-state value, diverges inversely with the distance from the critical point. The fact that this exponent is the same for all system parameters indicates a form of universality. The scaling behavior depends only on the structure of the underlying dynamical map and not on details such as the choice of initial state.
Repeated weak measurements with post-selection can produce abrupt changes in long-time meter observables, revealing weak values as tunable control parameters and enabling high-sensitivity detection, controlled manipulation of the meter state, and the exploration of critical-like dynamics even in small quantum systems.
@@ II Protocol presentation
This section presents the protocol that generates discontinuities and stability bifurcations and analyzes the corresponding Kraus operator. Several qubits serve as the “systems” in a standard weak measurement protocol with post-selection, while an $N$-level system acts as the meter, the system in which discontinuities are induced.
@@
The system and meter interact weakly through a unitary operator $\hat{U}$ which, assuming the interaction strength $g$ and the time $t$ are sufficiently small, can be expanded to first order in $gt$ as
@@
is the weak value of the observable $\hat{O}_{S}$. Accurately modeling the weak interaction and post-selection with the Kraus operator in Eq. 4 demands more than the condition $gt\ll 1$. It also requires that the following condition is satisfied [29]
$|gt\,O_{S,w}|\left(|\langle\hat{O}_{A}\rangle|+\Delta\hat{O}_{A}\right)\ll 1,$ (6)
where $\langle\hat{O}_{A}\rangle$ and $\Delta\hat{O}_{A}$ denote the expectation value and standard deviation of the ancilla observable $\hat{O}_{A}$ in its initial state.
@@
The weak measurement protocol to reach the discontinuity (Fig. 1), consisting of pre-selection, a weak unitary interaction, and post-selection, is applied to the meter, with the entire process represented by the Kraus operator $\hat{K}$. After each full application of the weak measurement protocol, the system is discarded and replaced with a new one for the next iteration (or reset to the initial state), while the meter is preserved throughout all $n$ applications.
@@ III Fixed points and their stability
$\hat{O}_{A}|a_{j}\rangle=o_{A,j}|a_{j}\rangle,\qquad\lambda_{j}\in\mathbb{R},$ (10)
where $j$ runs from 1 to $N$, the dimension of the Hilbert space of the operator $\hat{O}_{A}$, and $o_{A,j}$ are its eigenvalues.
@@
$|c_{j}^{(n)}|^{2}=\frac{|c_{j}^{(0)}|^{2}\,\left|1+2gt\,\text{Im}(O_{S,w})\,o_{A,j}\right|^{2n}}{\sum_{k}|c_{k}^{(0)}|^{2}\left|1+2gt\,\text{Im}(O_{S,w})\,o_{A,k}\right|^{2n}},$ (12)
@@
As $n\to\infty$, the system approaches a long-time limit characterized by a stable fixed point, which is the state toward which all other states converge, except for those corresponding to unstable fixed points. This stable state corresponds to the eigenvector of the Kraus operator associated with the eigenvalue of largest modulus, provided that the spectrum of the Kraus operator is non-degenerate. The remaining eigenvectors correspond to unstable fixed points.
@@
  • If $\text{Im}(O_{S,w})<0$, the stable state is the eigenvector with the largest eigenvalue $o_{A,+}$.
  • If $\text{Im}(O_{S,w})>0$, the stable state is the eigenvector with the smallest eigenvalue $o_{A,-}$.
@@ IV Qubit-qubit example
Figure 2: Three distinct discontinuities characterize the dependence of the expectation value of $\hat{\sigma}_{x}$ on the post-selected angle $\phi$, emphasizing the nontrivial structure that arises for different repetition numbers $n$. In this plot $\alpha=\frac{\pi}{3}$, $\theta=\frac{\pi}{4}$ and $\gamma=0.1$, while the initial meter state is along the $z$ axis.
@@
Figure 3: The imaginary part of the weak value vanishes when the absolute value of the two eigenvalues of the Kraus operator coincide. The plot shows the eigenvalues of the Kraus operator $\hat{K}$, $|\lambda_{+}|$ and $|\lambda_{-}|$, as functions of the post-selected angle $\phi$, together with their difference $|\lambda_{+}|-|\lambda_{-}|$ and the imaginary part of the weak value $\mathrm{Im}(O_{s,w})$. In this plot $\alpha=\frac{\pi}{3}$, $\theta=\frac{\pi}{4}$ and $\gamma=0.1$, while the initial meter state is along the $z$ axis.
As predicted analytically, the eigenvalues have equal absolute values whenever the imaginary part of the weak value vanishes (Fig. 3). The sum of the weak value weighted by the post-selection probability over all post-selected states corresponds to the expectation value of the operator. However, the sum of the imaginary part over all polar angles $\phi$ is nonzero, because one must also sum over all azimuthal angles $\alpha$ to account for all possible post-selected states.
@@
$r_{x,n}=\text{Tr}\left[\hat{\sigma}_{x}\hat{\rho}_{A}^{n}\right]=\frac{\kappa r_{x,n-1}+\beta+\zeta}{\kappa+\left(\beta+\zeta\right)r_{x,n-1}}$ (20)
$r_{y,n}=\text{Tr}\left[\hat{\sigma}_{y}\hat{\rho}_{A}^{n}\right]=\frac{\kappa r_{y,n-1}+ir_{z,n-1}\left(\beta-\zeta\right)}{\kappa+\left(\beta+\zeta\right)r_{x,n-1}}$
$r_{z,n}=\text{Tr}\left[\hat{\sigma}_{z}\hat{\rho}_{A}^{n}\right]=\frac{\kappa r_{z,n-1}-ir_{y,n-1}\left(\beta-\zeta\right)}{\kappa+\left(\beta+\zeta\right)r_{x,n-1}},$
where $\kappa=|c\left(\gamma\right)|^{2}+|d\left(\gamma\right)|^{2}$, $\beta=c\left(\gamma\right)d^{*}$, $\zeta=c^{*}\left(\gamma\right)d\left(\gamma\right)$.
@@
For values of $\phi$ such that $\phi<\frac{\pi}{2}$, two fixed points emerge, corresponding to the eigenvectors of $\hat{\sigma}_{x}$, the operator acting on the meter in Eq. 16. The fixed point at x=1x=-1 is stable, while the one at $x=1$ is unstable. Even if the state begins close to $x=1$, the dynamics will eventually drive it toward the stable fixed point at x=1x=-1 (Fig. 5). The dynamics will remain at $x=1$ only when the system is initially exactly at that point. Conversely, for $\frac{\pi}{2}<\phi<\pi$, the stability of the fixed points is reversed: $x=1$ becomes stable and x=1x=-1 unstable. After passing $\phi=\frac{\pi}{2}$, the stable and unstable fixed points effectively swap roles. Finally, in the near-unitary regime, when the dynamics are close to but not exactly unitary, the trajectories cycle on the Bloch sphere before eventually collapsing onto the stable fixed point (Fig. 6).
@@
The plot shown in Fig. 2 corresponds to the weak-measurement regime. However, as shown in Eq. 17, the dynamics can be conveniently analyzed for larger interaction strengths in this specific two-qubit case. For any value of $\gamma t$, the discontinuities persist in the three cases where the imaginary part of the weak value vanishes, namely $\phi=0$, $\phi=\frac{\pi}{2}$, and $\phi=\pi$. Nevertheless, the detailed behavior of the expectation value after a few applications of the Kraus operator depends on the interaction strength. As $\gamma t$ increases, the curve may initially deviate further from the zero axis during the first iterations of the protocol. For even stronger interactions, however, the behavior approaches that of the weak-measurement limit again, reflecting the underlying periodic dependence on $\gamma t$.
Moreover, the sign of the expectation value $\langle\hat{\sigma}_{x}\rangle$ in the intervals $\phi\in(0,\frac{\pi}{2})$ and $\phi\in(\frac{\pi}{2},\pi)$ changes at $\gamma t=\frac{\pi}{2}$, and subsequently alternates periodically with period $\frac{\pi}{2}$. At this point ($\gamma t=\frac{\pi}{2}$), the eigenvalue with the largest modulus changes, causing the previously unstable fixed point to become stable and vice versa. This exchange of stability gives rise to an additional discontinuity when $\gamma t$ is taken as the control parameter instead of $\phi$ (as shown in Fig. 7). For stronger interactions, the stability of the fixed points no longer depends solely on the sign of the imaginary part; rather, it becomes explicitly dependent on the interaction strength, leading to qualitative differences in stability across regimes. Fig. 7 shows the expectation value as a function of $\gamma t$ for a representative point within the first window of Fig. 2 $\left(\phi=\frac{\pi}{3}\right)$. As can be seen, discontinuities occur at intervals of $\frac{\pi}{2}$ in $\gamma t$. In contrast to the discontinuities observed when $\phi$ is used as the control parameter (Fig. 2), those appearing as a function of $\gamma t$ exhibit central symmetry (Fig. 7).
Figure 7: Discontinuities arise in the expectation value of $\hat{\sigma}_{x}$ as a function of $\gamma t$ at $\gamma t=m\frac{\pi}{2}$, where $m\in\mathbb{N}$, when the protocol is repeated a large number of times. In this plot, the parameters are fixed to $\alpha=\frac{\pi}{3}$, $\theta=\frac{\pi}{4}$, $\phi=\frac{\pi}{3}$, and the initial meter state is prepared along the z axis.